OpenStudy (anonymous):

MEDAL AND FAN!!!! PLZ SOLVE!!! A car travels down the highway at speeds between 50 miles per hour and 65 miles per hour. If the car travels for 4 hours, write an expression that represents the distance, d, the car has traveled.

OpenStudy (anonymous):

@pooja195

OpenStudy (misty1212):

what is \(4\times 50\)?

OpenStudy (anonymous):

200

OpenStudy (misty1212):

yeah, and how about \(4\times 65\)?

OpenStudy (anonymous):

260

OpenStudy (misty1212):

ok so the most it can go is 260 right? and the least is 200

OpenStudy (anonymous):

i think so???

OpenStudy (misty1212):

if the distance is D, then you could say \[200\leq D\leq 260\] meaning the distance has to be between those two numbers

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

how did u get it tho?

OpenStudy (anonymous):

Nvm its correct :) thank you